Investment analysis question 2.19

19. Investment Criteria. If you insulate your office for $1,000, you will save $100 a year in heating expenses. These savings will last forever a. What is the NPV of the investment when the cost of capital is 8 percent? 10 percent? b. What is the IRR of the investment? . What is the payback period on this investment?
Answer #1

a. What is the NPV of the investment when the cost of capital is 8%? 10%?

As the saving per year is last forever means perpetual

Therefore present value (PV) of these savings = annual saving / cost of capital

If cost of capital is 8%; then

PV of savings = $100/8% = $100/0.08 = $1,250

And net present value (NPV) = PV of savings – Initial investment

= $1,250 - $1,000 = $250

If cost of capital is 10%; then

PV of savings = $100/10% = $100/0.1 = $1,000

And net present value (NPV) = PV of savings – Initial investment

= $1,000 - $1,000 = $0

b. What is the IRR of the investment?

Internal Rate of Return (IRR) is the discount rate at which the Net Present Value (NPV) of an investment is zero and we know that at 10% discount rate NPV of investment is zero.

Therefore IRR is 10%

c. What is the payback period on this investment?

Payback period for an investment which has constant flow of annual income-

Payback period = Initial investment / Annual cash flow

= $1,000 / $100 = 10 years